OpenStudy (anonymous):

From points A and B, the distance between in which is 1020 mi. Two trains left simultaneously towards each other. The speed of one train was 10 mph greater than the speed of the other. In 5 hours the train had not met yet and were 170 mi apart. Find the speed of both trains.

hero (hero):

I would subtract 1020 - 170 to get the true distance they actually traveled first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Hero then wat would i do?

hero (hero):

Next you will add the distances together and set them equal to 850

hero (hero):

5x + 5(x + 10) = 850 5(x + x + 10) = 850 x + x + 10 = 850/5 2x + 10 = 170 2x = 170 - 10 2x = 160 x = 160/2 x = 80 x + 10 = 90

hero (hero):

Thus the speed of one train was 80 mph. The speed of the other was 90 mph

OpenStudy (anonymous):

thank you so much @Hero

hero (hero):

yw